Most listed Australian companies pay dividends twice per year, the 'interim' and 'final' dividends, which are roughly 6 months apart.

You are an equities analyst trying to value the company BHP. You decide to use the Dividend Discount Model (DDM) as a starting point, so you study BHP's dividend history and you find that BHP tends to pay the same interim and final dividend each year, and that both grow by the same rate.

You expect BHP will pay a $0.55 interim dividend in six months and a $0.55 final dividend in one year. You expect each to grow by 4% next year and forever, so the interim and final dividends next year will be $0.572 each, and so on in perpetuity.

Assume BHP's cost of equity is 8% pa. All rates are quoted as nominal effective rates. The dividends are nominal cash flows and the inflation rate is 2.5% pa.

What is the current price of a BHP share?
Answer #1

Initial Interim Dividend = $ 0.55 and Initial Yearly Dividend = $ 0.55

Effective Nominal Rate = r = 8%

Total Dividend Value at the end of Year 1 = 0.55 x (1.08)^(1/2) + 0.55 = $ 1.12158

Growth Rate of Dividends = g = 4 %

Therefore, Current Stock Price = 1.12158 / (r-g) = 1.12158 / (0.08-0.04) = $ 28.0394 ~ $ 28.039